OpenStudy (anonymous):

Which fraction is equivalent to 12%? A. B. C. D.

OpenStudy (anonymous):

A.12/25 B.48/25 C.3/25 D.1/8

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

If you want to find equivalent fractions to, say, 3/4, you can multiply the numerator and denominator by the same number.

OpenStudy (anonymous):

OK SO

OpenStudy (anonymous):

To find equivalent fractions, multiply the numerator AND denominator by the same nonzero whole number. This procedure is used to solve example 4. You can multiply the numerator and the denominator of a fraction by any nonzero whole number, as long as you multiply both by the same whole number!

OpenStudy (anonymous):

CAN U JUST GIVE ONLY ONE DRECT I NEED TO FINISH

OpenStudy (anonymous):

CAN U HELP

OpenStudy (anonymous):

12% = 0.12 = 12/100 = 3/25

OpenStudy (anonymous):

IF U DONTKNOW TELL ME SO I CAN GUESS

OpenStudy (anonymous):

THANKS